[Hallc_running] [New Logentry] RC Daily Update

billlee at jlab.org billlee at jlab.org
Thu Apr 11 11:05:02 EDT 2019

Logentry Text:
Dear Collaborators,
We had a mixed day of running.

Things started well yesterday, we completed running at Q2=5.5, low epsilon, theta=8.5 degree by 1pm yesterday. 1-3pm, Peter took over and performed SIDIS studies at high Q2, which required a lower SHMS momentum. Radcon swept the hall at 3pm, magnet experts went in to change the SHMS angle to 6.2 dgrees at 4pm, and the hall was back to beam permit at 5pm. We were set for the final setting of the K L/T at 50 uA.

Last night around 11pm, the SHMS led flow valve gave fault signal, and SHMS started to ramped down on its own. The interlock refuse to clear. After many communications with the experts, the dipole issue was not resolved, and  SHMS was down for the night. We switched the program to the HMS single arm optics running. The hall C experts (Steve Lassiter, Jerry, Andy) accessed the hall and power cycled the valve motor controller for the SHMS dipole at 8:30am this morning. The SHMS dipole was back online at 9 pm and completed ramping at 9:45am.  Currently, the accelerator is experiencing a PSS drop issue, and they are changing out the problematic power supply. Beam is expected around noon.

We will have a 3:30 meeting today (2nd floor of the counting house),

Many thanks


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3677919
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190411/9bcea701/attachment.html>

More information about the Hallc_running mailing list